# Unsocial Intelligence: a Pluralistic, Democratic, and Participatory Investigation of AGI Discourse

### v2

@@ 3.2.3. Human-like processes versus outcomes
Another divisive choice in conceptualizing AGI or human-level AI is whether to insist on mechanisms or processes that mimic human cognitive processes. For example, Morris et al. (2023) propose that most conceptions of AGI prioritize outcomes (“what an AGI can accomplish”) rather than human-like processes, because outcomes are more operationalizable. Excluding “processes” from the definition means that criteria based on the following are excluded: consciousness (Butlin et al., 2023; Lenharo, 2024; Searle, 1980; Summerfield, 2023; Smart, 2015), sentience (Schwitzgebel and Garza, 2015), and abilities modelled on human children (Turing, 2012; Nilsson, 2005; Gopnik, 2019, [n. d.]; Summerfield, 2023).
This choice of excluding anthropomorphic cognitive processes goes beyond purely conceptual questions about how to best define AGI. Discourse on AGI or human-level AI often takes a much more practical stance on this question: not about whether empirical assumptions should be in principle built into definitions, but rather about what directions are worth investing resources in on the path to AGI. [^3] Here is a possible epistemic argument for taking a stance on the exclusion of anthropomorphic processes from the concept of AGI. The question of whether or not implementing human-like processes is needed to achieve human-like outcomes is an empirical question. Settling open-ended empirical questions through definitions is bad epistemic practice. Definitions of AGI should, therefore, be agnostic on empirical questions such as the importance of human-like processes to achieve different human-like outcomes. We see this as amounting to providing a principled reason for not answering the question of whether AGI needs to involve anthropomorphic cognitive processes. Discourse on AGI often goes beyond this agnostic stance for practical reasons. This is the kind of practical choice that gets expressed in decisions about funding and resource allocation—Lenharo (2024) interviews a researcher claiming that “to his knowledge, there was not a single grant offer in 2023 to study the topic” of consciousness in AI.
Summerfield (2023) highlights the importance of Rich Sutton’s “The Bitter Lesson” in articulating this practical concern (Sutton, 2019). Sutton frames this as a choice between focusing on implementing human-like processes versus focusing on “general methods that leverage computation”. He frames the choices as practically exclusive in the sense that “[t]ime spent on one is time not spent on the other” Sutton (2019). The bitter lesson is that building human-like processes into AI may be psychologically satisfying to researchers, but tends to be much less effective than focusing on “general purpose method[s]” that scale with computation.
@@ 3.2.4. Generality
Although very widespread, the strong emphasis on “generality, adaptability and flexibility” (Goertzel, 2014) in machine intelligence is not universal among accounts of human-level machine intelligence. Bostrom (2014) is an example of someone who departs from the norm. Bostrom strongly rejects the requirement of human-like cognitive processes. He argues against assuming that achieving human-level intelligence or superintelligence requires “general” intelligence at all. Bostrom is agnostic on the structure of abilities needed to match or outperform the instrumental performance of human intelligence.
We argued earlier (Section 3.2.3) that assumptions about whether AGI should include human-like processes or not is value-laden because it is based on notions of “usefulness” or “effectiveness” that presuppose whom and what the assumptions are useful for. The same line of thought applies to arguments about whether generality of abilities is important. Assumptions about what practical outcomes are desirable are needed to make the case for or against the importance of generality. Bostrom’s line of argument presumes that replicating the instrumental performance of human intelligence is the goal—but there are other possible goals for AI. Similarly, others’ assumption that being polymathic is essential to AGI glosses over two points: whether this is a desirable goal for AI, and the real diversity of human abilities (not all humans are similarly polymathic). Here we see an analogy with critiques of g on the basis that it ignores human diversity (see Section 2.2.1).
@@ 3.2.5. Individualism
By contrast, accounts like Bostrom (2014) and Attard-Frost (2023) reject ontological individualism. Bostrom argues that when thinking about what it means for machines to match or outclass human intelligence, the relevant unit of comparison includes not just the intelligence of an individual human but also “the combined intellectual capability of all of humanity” at present (Bostrom, 2014). Specifically, he argues that collectives—such as “firms, work teas, gossip networks, advocacy groups, academic communities, countries, even humankind as a whole” (Bostrom, 2014)—can be understood as a mechanism for increasing intelligence. Bostrom also refers to collective intelligence as “collective intellectual problem-solving capacity”. We see this rejection of ontological individualism as related to the longer tradition of thinking about collective intelligence (Lévy, 1994, 2010; Engelbart, 1962; Baltzersen, 2021; Suran et al., 2021; Araya and Marber, 2023; Landemore and Elster, 2012; Anderson, 2006; Putnam, 2011; Dewey, 2011; Peters and Heraud, 2015). Baltzersen (2021) proposes thinking of collective intelligence as “collective problem solving”, both in large and small groups of people. [^6] Lévy (2010) similarly defines collective intelligence as “the capacity of human collectives to engage in intellectual cooperation in order to create, innovate and invent”. In Section 5, we come back to a strand of this tradition that frames democracy as a “precondition for the full application of intelligence to the solution of social problems” (Putnam, 2011).
@@ 3.2.6. Instrumentality Thesis
We view the assumption that intelligence is purely about instrumental rationality to be another juncture where values come into play. An alternate conception of intelligence could consider the ability to determine what goals are worth pursuing as a form of intelligence. Agents with this ability would be able to form independent views about whether the goals they are assigned by other agents are worthwhile, and if not, how or whether to resist their assignments. In short, the view that intelligence encompasses only instrumental rationality lays the ground for AI to develop agents that can optimize the goals of their designers, but that have no ability to question those goals. We can imagine instead a very different future of AI: where we have intelligent agents that are able to dissent (see Section 5). Which future is more desirable is a question of values: do we want intelligent machines that are completely subservient, or do we want machines that can collaborate with humans in determining what goals are worth pursuing?
@@ 3.2.7. Operationalizability/Measurability
Another dimension on which accounts of AGI differ is whether to require operationalizability or measurability for their definition. At one end of the spectrum, we have Morris et al. (2023), who frame their project around operationalizability and explicitly reject criteria that are not operationalizable. However, other discussants include definitions that are not measurable. Bostrom (2014), for example, makes frequent reference to the collective intelligence of all of humanity at present in his framework—a concept that seems difficult if not impossible to operationalize.
Operationalizability may appear value-neutral at first glance. However, scholars have argued that simply having a common language and shared standards is in itself value-laden. Commensuration (the act of establishing common standards to measure things) allows large institutions to coordinate actions and automate decision-making, but less powerful groups often argue that some things cannot be measured in order to resist actions that can be justified through “rational choices” (Espeland and Stevens, 1998). Social scientists have also recognized that commensuration has performative effects, creating a kind of path dependency: “as commensuration gets built into practical organizations of labor and resources, it becomes more taken for granted and more constitutive of what it measures” (Espeland and Stevens, 1998). This type of path dependency has been recognized in various domains of commensuration, such as standardized grades of grain quality (Porter, 1996) and university rankings (Fowles et al., 2016). Previous work on ML benchmarking has also highlighted how benchmarking creates performativity and path dependency (Blili-Hamelin, Borhane and Hancox-Li, 2023).
@@ 3.2.8. Choice of tasks/benchmarks
One cross-cutting dimension of difference across definitions of AGI is the tasks or benchmarks that they include in their criteria for AGI. Some of the previously mentioned dimensions intersect with this—for example, if embodiment is part of your AGI definition, you would include some embodied tasks in your definition. In contemporary AI research with its plethora of tasks and benchmarks, researchers sometimes make choices about which tasks they are emphasizing when estimating progress towards AGI.
Two prominent AI researchers recently argued that “[t]he most important parts of AGI have already been achieved by the current generation of advanced AI large language models” (Agüera y Arcas and Norvig, 2023). While these researchers define AGI roughly as instructable systems able to operate over a wide variety of topics, tasks, modalities and languages, the fact that the strengths of current LLMs are considered to already cover “the most important parts of AGI” implies that the things that LLMs cannot do are the less important parts of AGI. These LLM weaknesses include (as they acknowledge themselves) arithmetic and adhering to facts. However, they do not explain why these areas of reasoning are devalued relative to LLM strengths like translation or synthesizing code. Other researchers, in contrast, believe that arithmetical and logical reasoning are necessary components of AGI (Marcus, 2023). Agüera y Arcas and Norvig (2023)’s assertion that what LLMs are good at constitute the most important aspects of AGI is value-laden. Choices of tasks and benchmarks are a key site of value-laden design choices (Blili-Hamelin, Borhane and Hancox-Li, 2023; Bommasani, 2022). Deciding to focus on LLMs over deductive reasoning systems has ethical implications because these systems will change our societies in different ways.  [^8] For another critique of the reliance on benchmarks in relation to AGI, see Summerfield (2023).
@@ 3.2.9. Importing g into AI
$3.$ Are deflationary views truly deflationary?
@@
Furthermore, there are insights to be drawn from Morris et al. (2023)’s analogy to the Society of Automotive Engineers’ (SAE) framework of levels of automation for automobiles (International, 2021). The intention of the analogy was to justify a similar move in the AI space, with the implied utility of a common language and shared standards. However, critical work on the SAE framework has argued that what appears to be merely a descriptive, technical definition with the SAE’s stated intentions of “simplifying communication” and “providing clarity” in fact has normative assumptions and implications (Hopkins and Schwanen, 2021). Implications teased out in (Hopkins and Schwanen, 2021) include: promoting “homogeneity in mobility futures”, specifically a homogeneous vision of automation as the future of mobility; reproducing a “dominant discourse of an expert-led, technologically-centred vision of mobility futures”; and removing “obstacles and impediments to the successful (and timely) development of the automated vehicle niche.” Similarly, we may ask if shared standards for AGI promote homogeneity in AI futures and reproduce an expert-led dominant discourse. Certainly, the standards currently being developed are expert-led and lack input from people who are more likely to suffer negative impacts in AI—an issue that we discuss further in Section 4.3.
4. Towards contextualized, democratic, and social intelligence
We now outline alternative visions for values worth centering in imagining transformative machine intelligence. These views embrace the value-laden nature of intelligence instead of side-stepping or ignoring it. Broadly speaking, the following proposals take seriously the role of physical and social contexts in definitions of AGI, in contrast to the views outlined in Section 3. We selected these views because they embody the values of contextualism, epistemic justice, inclusiveness, and democracy, which we consider vital for visions of the future of AI worth pursuing.
@@ 4.1. Contextual intelligence
Attard-Frost’s account may have some practical drawbacks—perhaps it is harder to operationalize, or to compare systems that operate in different contexts. Thus, whether or not we choose to adopt something like their account is itself dependent on our values and goals: their goal of queering intelligence is itself a normative one, and may conflict with other normative goals.
@@ 4.2. Inductive risk and social values
Another helpful lens on how to define AGI is that of inductive risk. In the values in science literature, inductive risk is often used to justify having different standards of evidence for accepting scientific claims. The idea is that when the potential social costs of accepting/denying a scientific claim are high, social and political values should influence standards of evidence (Douglas, 2000; Steel, 2010; Blili-Hamelin, Borhane and Hancox-Li, 2023).
We can adapt the argument from inductive risk to the case of defining AGI, but reframe it in terms of standards of evidence for accepting definition(s) of AGI. [^12] We frame it this way because it is unclear whether definitions of AGI are “scientific claims” as traditionally construed by philosophers of science. We thank Ravit Dotan for proposing this point. As Section 3 has illustrated, current AGI definitions make many choices that are unjustified by epistemic values, or are based on implicit assumptions about what is socially valuable. Adapting the argument from inductive risk to AGI, we would say: due to the significant social impacts that could arise from having definitions of intelligence that value certain social goals/tasks/beings over others, social and political values should influence how we define AGI.
4.3. Epistemic justice for defining AGI
One aspect of many definitions of AGI that we have analyzed here is that they mostly come from actors who have relatively more power and influence over the future of AI. At the same time, there is a dearth of voices in the conversation from people who are more likely to be harmed by deployed AI systems. Advocates of AGI and human-level AI imagine this technology as impacting almost everyone. Taking that ambition seriously requires processes that give a meaningful say to the communities who would be impacted by the technology. Dreams of building transformative technologies should come hand in hand with participatory, inclusive, and—as we argue below— democratic decision-making processes.
A positive vision for defining AGI (or, taking a step back, even deciding if it should be measured) would learn from the participatory ML and epistemic justice literature. We would love to see a vision for AGI—or better, transformative forms of machine intelligence—that is constructed through participatory methods (Birhane et al., 2022a; Delgado et al., 2023), while still being aware that the products of these methods have their limitations (Sloane et al., 2020). Needless to say, these participatory methods should take care to include the perspectives of not just those who design or fund AI, but those who will be impacted by it in other ways.
@@ 4.3. Epistemic justice for defining AGI
5. Conclusion: democratic intelligence
We have highlighted value-laden choices in discourse about human-level AI and AGI. We surfaced alternative paths that provide more contextual, participatory, and epistemically just perspectives on the topic. We suggested that the question of what forms of machine intelligence are worth pursuing calls for a plurality of contested and value-laden perspectives. We now conclude by highlighting perspectives that provide positive rejoinders for placing dissent, deliberation, and political legitimacy at the center of conversations about machine intelligence and transformative technologies.

### v3

@@ 3.2.3. Human-like processes versus outcomes
Another divisive choice in conceptualizing AGI or human-level AI is whether to insist on mechanisms or processes that mimic human cognitive processes. For example, Morris et al. (2023) want to prioritize outcomes (“what an AGI can accomplish”) over human-like processes because outcomes are more operationalizable. Excluding “processes” from the definition means that criteria based on the following are excluded: consciousness (Butlin et al., 2023; Lenharo, 2024; Searle, 1980; Summerfield, 2023; Smart, 2015), sentience (Schwitzgebel and Garza, 2015), and abilities modeled on human children (Turing, 2012; Nilsson, 2005; Gopnik, 2019, 2015; Summerfield, 2023).
This choice of excluding anthropomorphic cognitive processes goes beyond purely conceptual questions about how to best define AGI. [^3] Here is a possible epistemic argument for taking a stance on the exclusion of anthropomorphic processes from the concept of AGI. The question of whether or not implementing human-like processes is needed to achieve human-like outcomes is an empirical question. Settling open-ended empirical questions through definitions is bad epistemic practice. Definitions of AGI should, therefore, be agnostic on empirical questions such as the importance of human-like processes to achieve different human-like outcomes. We see this as amounting to providing a principled reason for not answering the question of whether AGI needs to involve anthropomorphic cognitive processes. Discourse on AGI often goes beyond this agnostic stance for practical reasons. It also influences the types of research projects that attract AI investment, which has ethical consequences in terms of what the AI we build (and its attendant social consequences) looks like. Lenharo (2024) interviews a researcher claiming that “to his knowledge, there was not a single grant offer in 2023 to study the topic” of consciousness in AI.
Rich Sutton’s “The Bitter Lesson” articulates this practical concern about maximizing return on investment (Sutton, 2019). Sutton frames this as a choice between focusing on implementing human-like processes versus focusing on “general methods that leverage computation”. He frames the choices as practically exclusive in the sense that “[t]ime spent on one is time not spent on the other” Sutton (2019). The bitter lesson is that building human-like processes into AI may be psychologically satisfying to researchers, but tends to be much less effective than focusing on “general purpose method[s]” that scale with computation.
@@ 3.2.4. Generality
Although widespread, the strong emphasis on “generality, adaptability and flexibility” (Goertzel, 2014) in machine intelligence is not universal among accounts of human-level machine intelligence. For instance, Bostrom (2014) departs from this norm by denying that achieving human-level intelligence or superintelligence requires “general” intelligence at all. Bostrom sees this as part of his strong rejection of the requirement for human-like cognitive processes. He instead remains agnostic about the structure of abilities needed to match or outclass the instrumental performance of human intelligence.
We argued earlier (Section 3.2.3) that assumptions about whether AGI should include human-like processes or not is value-laden because it is based on notions of “usefulness” or “effectiveness” that presuppose whom and what the assumptions are useful for. The same line of thought applies to arguments about whether the generality of abilities is important. Assumptions about what practical outcomes are desirable are needed to make the case for or against the importance of generality. Bostrom’s line of argument presumes that replicating the instrumental performance of human intelligence is the goal—but there are other possible goals for AI. Similarly, others’ assumption that being polymathic is essential to AGI glosses over two points: whether this is a desirable goal for AI, and the real diversity of human abilities (not all humans are similarly polymathic). Here we see an analogy with critiques of g on the basis that it ignores human diversity (see Section 2.1.1).
@@ 3.2.5. Individualism
By contrast, accounts like Bostrom (2014) and Attard-Frost (2023) reject ontological individualism. Bostrom argues that when thinking about what it means for machines to match or outclass human intelligence, the relevant unit of comparison includes not just the intelligence of an individual human but also “the combined intellectual capability of all of humanity” at present (Bostrom, 2014). Specifically, he argues that collectives—such as “firms, work teas, gossip networks, advocacy groups, academic communities, countries, even humankind as a whole” (Bostrom, 2014)—can be understood as a mechanism for increasing intelligence. Bostrom also refers to collective intelligence as “collective intellectual problem-solving capacity”. This rejection of ontological individualism is related to the longer tradition of thinking about collective intelligence (Lévy, 1994, 2010; Engelbart, 1962; Baltzersen, 2021; Suran et al., 2021; Araya and Marber, 2023; Landemore and Elster, 2012; Anderson, 2006; Putnam, 2011; Dewey, 2011; Peters and Heraud, 2015). Baltzersen (2021) proposes thinking of collective intelligence as “collective problem solving”, both in large and small groups of people. [^6] Lévy (2010) similarly defines collective intelligence as “the capacity of human collectives to engage in intellectual cooperation in order to create, innovate and invent”. In Section 5, we come back to a strand of this tradition that frames democracy as a “precondition for the full application of intelligence to the solution of social problems” (Putnam, 2011).
@@ 3.2.6. Instrumentality Thesis
The assumption that intelligence is purely about instrumental rationality is another juncture where values come into play. An alternate conception of intelligence could consider the ability to determine what goals are worth pursuing as a form of intelligence. Agents with this ability would be able to form independent views about whether the goals they are assigned by other agents are worthwhile, and if not, how or whether to resist their assignments. In short, the view that intelligence encompasses only instrumental rationality lays the ground for AI to develop agents that can optimize the goals of their designers, but that have no ability to question those goals. We can imagine instead a very different future of AI: where we have intelligent agents that are able to dissent about final goals (see Section 5). Which future is more desirable is a question of values: do we want intelligent machines that are completely subservient, or do we want machines that can collaborate with humans in determining what goals are worth pursuing?
@@ 3.2.7. Operationalizability/Measurability
Another dimension on which accounts of AGI differ is whether to require operationalizability or measurability for their definition. At one end of the spectrum, we have Morris et al. (2023), who frame their project around operationalizability and explicitly reject criteria that are not operationalizable. However, other discussants include definitions that are not measurable. Bostrom (2014), for example, makes frequent reference to the collective intelligence of all of humanity at present in his framework—a concept that seems difficult, if not impossible, to operationalize.
Operationalizability may appear value-neutral at first glance. However, scholars have argued that simply having a common language and shared standards is in itself value-laden. Commensuration (the act of establishing common standards to measure things) allows large institutions to coordinate actions and automate decision-making, but less powerful groups often argue that some things cannot be measured in order to resist actions that can be justified through “rational choices” (Espeland and Stevens, 1998). Social scientists have also recognized that commensuration has performative effects, creating a kind of path dependency: “as commensuration gets built into practical organizations of labor and resources, it becomes more taken for granted and more constitutive of what it measures” (Espeland and Stevens, 1998). This type of path dependency has been recognized in various domains of commensuration, such as standardized grades of grain quality (Porter, 1996) and university rankings (Fowles et al., 2016). Previous work on ML benchmarking has also highlighted how benchmarking creates performativity and path dependency (Dehghani et al., 2021; Blili-Hamelin and Hancox-Li, 2023).
@@ 3.2.8. Choice of tasks/benchmarks
One cross-cutting dimension of difference across definitions of AGI is the tasks or benchmarks that they include in their criteria for AGI. Some of the previously mentioned dimensions intersect with this—for example, if embodiment is part of your AGI definition, you would include some embodied tasks in your definition. In contemporary AI research, with its plethora of tasks and benchmarks, researchers sometimes make choices about which tasks they emphasize when estimating progress toward AGI.
Two prominent AI researchers recently argued that “[t]he most important parts of AGI have already been achieved by the current generation of advanced AI large language models” (Agüera y Arcas and Norvig, 2023). While these researchers define AGI roughly as instructable systems able to operate over a wide variety of topics, tasks, modalities and languages, the fact that the strengths of current LLMs are considered to already cover “the most important parts of AGI” implies that the things that LLMs cannot do are the less important parts of AGI. These LLM weaknesses include (as they acknowledge themselves) arithmetic and adhering to facts. However, they do not explain why these areas of reasoning are devalued relative to LLM strengths like translation or synthesizing code. Other researchers, in contrast, believe that arithmetical and logical reasoning are necessary components of AGI (Marcus, 2023). Agüera y Arcas and Norvig (2023)’s assertion that the most important aspects of AGI are achieved by current-day LLMs is value-laden. Choices of tasks and benchmarks are a key site of value-laden design choices (Blili-Hamelin and Hancox-Li, 2023; Bommasani, 2022). Deciding to focus on LLMs over deductive reasoning systems has ethical implications because these systems will change our societies in different ways.  [^8] For another critique of the reliance on benchmarks in relation to AGI, see Summerfield (2023).
@@ 3.2.9. Importing g into AI
3.3. Are deflationary views truly deflationary?
@@
Furthermore, there are insights to be drawn from Morris et al. (2023)’s analogy to the Society of Automotive Engineers’ (SAE) framework of levels of automation for automobiles (SAE International, 2021). The intention of the analogy was to justify a similar move in the AI space, with the implied utility of a common language and shared standards. However, critical work on the SAE framework has argued that what appears to be merely a descriptive, technical definition with the SAE’s stated intentions of “simplifying communication” and “providing clarity” in fact has normative assumptions and implications (Hopkins and Schwanen, 2021). Implications teased out by Hopkins and Schwanen include: promoting “homogeneity in mobility futures”, specifically a homogeneous vision of automation as the future of mobility; reproducing a “dominant discourse of an expert-led, technologically-centred vision of mobility futures”; and removing “obstacles and impediments to the successful (and timely) development of the automated vehicle niche.” Similarly, we may ask if shared standards for AGI promote homogeneity in AI futures and reproduce an expert-led dominant discourse. Certainly, the standards currently being developed are expert-led and lack input from people who are more likely to suffer negative impacts in AI—an issue that we discuss further in Section 4.3.
4. Towards contextualized, politically legitimate, and social intelligence
We now outline alternative visions for values worth centering in imagining future forms of machine intelligence. These views embrace the value-laden nature of intelligence instead of side-stepping it. Broadly speaking, the following proposals take seriously the role of physical and social contexts in definitions of machine intelligence, in contrast to the views outlined in Section 3. We selected these views because they embody the values of contextualism, epistemic justice, inclusiveness, and democracy, which we consider vital for visions of the future of AI worth pursuing.
@@ 4.1. Contextual intelligence
Attard-Frost’s account may have some practical drawbacks—perhaps it is harder to operationalize, or to compare systems that operate in different contexts. Thus, whether or not we choose to adopt something like their account is itself dependent on our values and goals: their goal of queering intelligence is itself a normative one, and it may conflict with other normative goals.
@@ 4.2. Inductive risk and social values
Another helpful lens on how to define future forms of machine intelligence is that of inductive risk. In the values in science literature, inductive risk is often used to justify having different standards of evidence for accepting scientific claims. The idea is that when the potential social costs of accepting/denying a scientific claim are high, social and political values should influence standards of evidence (Douglas, 2000; Steel, 2010; Blili-Hamelin and Hancox-Li, 2023).
We can adapt the argument from inductive risk to the case of defining AGI by reframing it in terms of standards of evidence for accepting definition(s) of AGI. [^12] We frame it this way because it is unclear whether definitions of AGI are “scientific claims” as traditionally construed by philosophers of science. We thank Ravit Dotan for proposing this point. As Section 3 has illustrated, current AGI definitions make many choices that are unjustified by epistemic values, or are based on implicit assumptions about what is socially valuable. Adapting the argument from inductive risk to AGI, we would say: due to the significant social impacts that could arise from having definitions of intelligence that value certain social goals/tasks/beings over others, social and political values should influence how we define AGI.
4.3. Epistemic justice for defining future forms of AI
One aspect of many definitions of AGI that we have analyzed here is that they mostly come from actors who have relatively more power and influence over the future of AI. At the same time, there is a dearth of voices in the conversation from people who are more likely to be harmed by deployed AI systems. Advocates of AGI and human-level AI imagine this technology as impacting almost everyone. Taking that ambition seriously requires processes that give a meaningful say to the communities who would be impacted by the technology. Dreams of future technologies should come hand in hand with participatory, inclusive, and—as we argue below— politically legitimate decision-making processes.
A positive vision for defining future forms of AI (or, taking a step back, even deciding if it should be pursued) would learn from the participatory ML and epistemic justice literature. We would love to see a vision for future forms of machine intelligence that is constructed through participatory methods (Birhane et al., 2022a; Delgado et al., 2023; Young et al., 2024), while still being aware that the products of these methods have their limitations (Sloane et al., 2020). Needless to say, these participatory methods should take care to include the perspectives of not just those who design or fund AI, but those who will be impacted by it in other ways.
@@ 4.3. Epistemic justice for defining future forms of AI
Notably, by suggesting more participatory methods, we do not mean that every issue surrounding future forms of machine intelligence should be subject to a population-wide referendum. As the IPCC case illustrates, specialist expertise is indispensable to making good decisions on these issues. Luckily, modern democracies have several methods of ensuring that citizens have a voice in state decisions through mechanisms other than referenda. For example, representative democracy, where citizens elect representatives who make decisions on behalf of them, mixes elements of expert-led decision making with citizen participation. Similarly, polities have also experimented with citizen assemblies on important issues (Devaney et al., 2020; Huang and Tu, 2017). These are small forums where a selected set of “ordinary” citizens debate issues as a mechanism for influencing public opinion and political decision-makers.
5. Conclusion: politically legitimate intelligence
Donna Haraway’s “A Manifesto for Cyborgs” argues that despite its origin in “racist, male-dominant capitalism” and “militarism”, the figure of the human-machine hybrid can be repurposed towards resisting oppressive social categories in imagining the future (Haraway, 1987; Forlano and Glabau, 2024). Imagining “altering bodily functions” to allow humans to survive in space (Clynes and Kline, 1960) opens the door to Ashley Shew’s interrogation of how technology reinforces ableism:“Everyone in space will be disabled” (Shew, 2018; Forlano and Glabau, 2024). The scale, power, and scope—or perhaps lack of scope, as Gebru and Torres argue (Gebru and Torres, 2024; Burrell and Metcalf, 2024)—of dreams of AGI should raise serious doubt about whether the concept is worth similarly subverting. Whether through subverting or discarding AGI, we nevertheless believe in the need for more contextual, participatory, and democratic approaches to imagining future forms of intelligence and machines.
To that end, we have taxonomized the discordant value-laden choices of AGI discourse. We surfaced alternative paths that provide more contextual, participatory, and epistemically just perspectives on imagining future forms of AI. We suggested that the question of what forms of machine intelligence are worth pursuing calls for a plurality of contested and value-laden perspectives. We now conclude by highlighting perspectives that provide positive rejoinders for placing dissent, deliberation, and political legitimacy at the center of conversations about intelligence and future technologies.